OpenStudy (anonymous):

-9√28a^2 * 1/3√63a step by step needed. I just don't understand it.

OpenStudy (solomonzelman):

\[-9\sqrt{28a^2}~~\times~~\frac{1}{3}\sqrt{63a}\]Lets take one thing at a time.

OpenStudy (solomonzelman):

\[-9\sqrt{28a^2}~~~~~->~~~~~-9 \times \sqrt{7 \times 4 \times a^2}\]

OpenStudy (solomonzelman):

So far so good?

OpenStudy (anonymous):

So what is the - > sign for?

OpenStudy (solomonzelman):

that means, "turns into"

OpenStudy (anonymous):

OH. okay I get it :)

OpenStudy (solomonzelman):

\[~~->~~~~~~-9 \times \sqrt{7} \times \sqrt{4} \times \sqrt{a^2}\] (say something like OK, or anything, so that I know you are following every time.)

OpenStudy (anonymous):

ok :)

OpenStudy (solomonzelman):

I got disconnected, sorry.

OpenStudy (anonymous):

That's okay :) It's happened to me a few times ahah

OpenStudy (solomonzelman):

\[-9 \times \sqrt{7} \times \color{red} {\sqrt{4}} \times \color{red} { \sqrt{a^2} }\]\[-9 \times \sqrt{7} \times \color{red} {\sqrt{2^2}} \times \color{red} { \sqrt{a^2} }\]\[-9 \times \sqrt{7} \times \color{red} {2} \times \color{red} { a }\]\[-18a \times \sqrt{7}\]

OpenStudy (solomonzelman):

So far?

OpenStudy (anonymous):

Great :)

OpenStudy (solomonzelman):

\[-18a \sqrt{7}\]is the first part, lets work on the second part.

OpenStudy (solomonzelman):

\[ \frac{1}{3}\sqrt{63a}\]

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

\[\frac{1}{3} \times \sqrt{9 \times 7 \times a}\]\[\frac{1}{3} \times \sqrt{9} \times \sqrt{7a}\]\[\frac{1}{3} \times \sqrt{3^2}\times \sqrt{7a}\]\[\frac{1}{3} \times 3 \times \sqrt{7a}\]\[1 \times \sqrt{7a}\]\[\sqrt{7a}\]

OpenStudy (anonymous):

okay :)

OpenStudy (solomonzelman):

So you have 2 parts \[\color{blue} {-18a \sqrt{7} } \times \color{red} {\sqrt{7a } }\]first part is in blue, the second is in red.

OpenStudy (solomonzelman):

\[\color{blue} {-18a \sqrt{7} } \times \color{red} {\sqrt{7 \times a} }\]\[\color{blue} {-18a \sqrt{7} } \times \color{red} {\sqrt{7} } \times \color{red} {\sqrt{a} }\]combine the red and blue sevens,

OpenStudy (anonymous):

okay :)

OpenStudy (solomonzelman):

My connection again.

OpenStudy (solomonzelman):

\[-18a \times \sqrt{7} \times \sqrt{7} \times \sqrt{a}\]\[-18a \times \sqrt{7^2} \times \sqrt{a}\]\[-18a \times 7 \times \sqrt{a}\]\[-126a \times \sqrt{a}\]\[-126a \sqrt{a}\]

OpenStudy (anonymous):

I was right! Thank you so much

OpenStudy (solomonzelman):

Anytime!